# HuggingGraph: Understanding the Supply Chain of LLM Ecosystem

### v1

@@ 4.4. RQ #4: Supply Chain between LLM Models and Datasets
Table 5. Top-10 models sorted by the # of datasets used for training.
Model
Model Type
# of dataset used-
for training
DeBERTa-ST-AllLayers-v3.1 Fine tune 116
DeBERTa-ST-AllLayers-v3.1bis Adapters 116
static-similarity-mrl-mul-v1 Fine tune 108
static-similarity-mrl-multilingual Fine tune 108
ModernBERT-base-embed Fine tune 88
Llama-3.2-3B-Instruct Fine tune 87
Llama-3.2-3B-Instruct-GGUF Quantization 87
DavidLanz-3.2-3B-Instruct Fine tune 87
static-retrieval-mrl-en-v1 Fine tune 79
XLMRoBERTaM3-CustomPoolin Fine tune 72
4.5. RQ #5: Temporal Update and Tracking
This research question aims to focus on the value of tracking fine-grained, time-based changes in the LLM supply chain. Our $\Delta$-based update mechanism (discussed in Section 3.4), produces a daily snapshot of the HuggingGraph, capturing precisely how many nodes and edges are added each day, how that translates into net growth, and what the busiest single day looks like. To better understand the implications of this tracking, we adopt both qualitative and quantitative perspectives on the evolving LLM ecosystem.
Fine-grained temporal insights into the LLM ecosystem. We would like to understand how the LLM supply chain evolves on a day-to-day basis. Hence, we adopt this fine-grained temporal qualitative approach that highlights fluctuations in model and dataset activity. This strategy allows us to uncover ecosystem-level dynamics such as release bursts, contributor behavior, and category-specific trends over time.
Figure 5 visualizes the daily $\Delta$-based growth of five key node categories (base, fine-tuned, adapters, quantized variants models, and datasets) from June 25 to July 15, 2025.
We observe that, (i) the fine-tuned models dominate the daily activity, averaging over 1,700 new entries per day, followed by consistent contributions from adapters ($\sim$750/day) and datasets ($\sim$700/day). Clear spikes on specific days—e.g., July 7 and July 9—coincide with major releases such as the Mistral-Fusion-v3 fine-tuning wave and dataset refreshes like HFTime2025-News. (ii) Furthermore, adapter uploads peaked at 887 on June 28, while quantized variants reached 290 on July 9, driven by releases such as QWin-GGUF-7B. These patterns demonstrate HuggingGraph’s ability to capture evolving supply chain dynamics at a fine-grained, time-sensitive resolution, revealing ecosystem bursts and contributor behaviors in real time.
Quantifying daily incremental graph growth. To complement our qualitative insights with measurable evidence, we adopt a quantitative approach that captures the scale and efficiency of daily changes to the LLM supply chain. This allows us to evaluate the performance of our $\Delta$-based update mechanism, ensuring support for large-scale, time-sensitive graph maintenance in a computationally efficient and auditable manner.
Table 6 provides a detailed breakdown of daily additions, net changes, and peak updates across node and edge types from Jun. 25 to Jul. 15, 2025. (i) On average, Hugging Face added 126 new fine-tuned models per day (peaking at 2,090), followed by adapters and datasets with 41 and 39 daily insertions, respectively. Quantized and merged models contributed 16 and 1 additions per day, respectively. In total, 2,526 fine-tuned models and 785 datasets were added during this period. (ii) Each day also saw approximately 23 $\pm$ 6 new dependency edges—linking models to their parent checkpoints or training datasets—consistent. By comparing successive graph versions, our update mechanism applies only the incremental changes ($\Delta_{t}$) rather than rebuilding the full graph—keeping updates both scalable and auditable in practice.
@@ 4.5. RQ #5: Temporal Update and Tracking
Figure 5. The temporal update for the datasets and different models on Hugging Face from June 25 to July 15, 2025.
Finding #6: The $\Delta$-based update mechanism offers clear visibility into the LLM ecosystem’s rapid and continuous evolution, with daily growth primarily driven by fine-tuned, adapter, and quantized models. It enables efficient, audit-friendly tracking by applying only incremental changes, eliminating the need to rebuild the entire graph. These capabilities ensure that HuggingGraph remains both scalable and precise in capturing the dynamic nature of the LLM supply chain.
@@ 5. Related Work
LLM supply chain perspectives in AI. LLMs mark the revolutionary era in AI, with the edge passing through an exponential growth phase. A recent paper presents LLMs key components, including model infrastructure, lifecycle, and downstream applications (Wang et al., 2024). Another work showed that reusing these models has become so widespread, which encourages the sharing and adaptation of foundation models on a larger scale (Jiang et al., 2023). An open-source AI ecosystem, such as Hugging Face, hosts a broad range of LLMs and datasets and, therefore, plays an important role in democratizing AI technologies (Riva et al., 2024). The foundation models are key ingredients in such an ecosystem (Xu et al., 2024b). They can capture vast amounts of information from various datasets, allowing for effective task-specific variants like fine-tuning (Han et al., 2021). This trend underscores the significant impact of technological advancements on the democratization and innovation in AI (Cupać et al., 2024).
Table 6. Daily $\Delta$ statistics for HuggingGraph updates (June 25–July 15, 2025), showing average additions, net changes, and peak values by node and edge type.
Category
New-
per day
Net Change-
per day
Peak-
in month
Datasets 39 $\pm$ 62 +785 849
Base model 2 $\pm$ 5 +57 42
Fine-tune 126 $\pm$ 168 +2526 2090
Adapter 41 $\pm$ 64 +828 887
Quantized 16 $\pm$ 19 +321 290
Merge 1 $\pm$ 2 +34 13
Dependency edges 23 $\pm$ 6 +189 289
@@
Relationship analysis between LLM models and datasets. A recent study investigates the practical adaptation of foundation models to specific domains and tasks. Multitask fine-tuning has demonstrated the potential to enhance performance on target tasks with scarce labels (Xu et al., 2024a). In plant phenotyping, adapting vision foundation models by techniques like adapter tuning and decoder tuning has shown results comparable to those of leading task-specific models (Chen et al., 2023). The Quadapter technique for language models tackles quantization difficulties by incorporating learnable parameters that scale activations channel-wise, mitigating overfitting during quantization-aware training (Park et al., 2022). The PathoTune framework in pathology imaging utilizes multi-modal prompt tuning to reconcile domain discrepancies among foundation models, tasks, and specific instances, exhibiting enhanced performance compared to single-modality methods and facilitating the direct adaptation of natural visual foundation models to pathological tasks (Lu et al., 2024a). These findings underscore the efficacy of various adaptation approaches in improving foundation model performance across multiple areas.
@@
From the above, we can see that models heavily depend on each other for fine-tuning, adaptation, or quantization. Building on these insights, two other works have indicated that serious security vulnerabilities may encumber the LLM supply chain to a few developers (Kathikar et al., 2023; Osborne et al., 2024). These factors may affect the diversity of innovation. The authors relate several challenges arising during the software engineering, security, and privacy of different relations and components involved in model creation and deployment (Huang et al., 2023; Wang et al., 2024). Moreover, the vulnerabilities can be transferable from one model to another model during the fine-tuning process (Hajipour et al., 2024). Therefore, serious security challenges are also posed because the vulnerabilities of the repositories threaten the whole ecosystem (Kathikar et al., 2023).
Comparison with HuggingGraph. Our work is different from prior works in three aspects. First, the motivation for this work stems from the transferable property (Hajipour et al., 2024) of one model to others and during the training process with the datasets. Because of this, our analysis is mainly about the hierarchical relationship between the models, which model is used to fine-tune, adapt, or quantize other models, and which datasets are used to train the models. As it is challenging to get consistent hierarchical relational information, we made a supply chain graph from the metadata using APIs we got from the models and datasets. Second, as our analysis is one of the premier tasks in constructing a supply chain graph, for analyzing the relationship, we focus on the different graph properties like connectivity and WCC analysis. Third, we use forward and backward supply chain analysis to examine the built graph and learn more about the hierarchies of the models and datasets during fine-tuning, adaptation, quantization, and training. This is different from previous methods, which mostly focused on training the models. This would facilitate determining and identifying the relationship between the model and the dataset. More importantly, we conclude with multiple interesting findings, which we believe could provide important insights to the LLM community.
6. Discussion
Use case #1: Tracing lineage and dependencies in the LLM supply chain. In the LLM ecosystem, models are frequently built upon others through fine-tuning, adapter training, or quantization, forming complex chains of dependencies. However, when these relationships are not explicitly visible, it becomes difficult to verify where a model comes from, whether it inherits bias from upstream datasets, or if it complies with licensing constraints. HuggingGraph can be used to address this challenge by constructing the supply chain of models and datasets, uncovering both direct and derived dependencies, even when they are not formally documented. For example, it can trace how the model Meta-llama indirectly relies on a dataset like Wikimedia via Awesome-Chatgpt-prompt (Figure 2). This transparency supports developers, auditors, and policymakers in validating provenance, detecting risks, and enabling responsible, trustworthy AI deployment.
Use case #2: Identifying critical nodes and structural vulnerabilities. In the LLM ecosystem, certain models (e.g., gemma-2b) and datasets (e.g., The Pile) are reused so frequently that they become critical structural hubs, where failure or removal of them could disrupt numerous downstream dependencies. These hidden single points of failure are difficult to detect without a comprehensive view of resource interconnections. HuggingGraph can be used to address this by modeling the supply chain as a graph and analyzing node connectivity to surface highly reused models and datasets with significant inbound or outbound links. This visibility enables maintainers to safeguard vital assets and helps developers mitigate the risk of overreliance on fragile or under-maintained components, fostering more robust and secure model development pipelines.
@@ 7. Conclusion
This project studies the relationships between models and datasets in the LLM ecosystem, which are the central parts of the LLM supply chain. First, we design a methodology to systematically collect the supply chain information of LLMs. With the collected information, we design a new graph to model their relationships, which is a large directed heterogeneous graph, having 397,376 nodes and 453,469 edges. From there, we perform different types of analysis and make multiple interesting findings.

### v2

@@ 4.5. RQ #5: Supply Chain between Models and Datasets
In addition, models like static-similarity-mrl-mul-v1 and static-similarity-mrl-multilingual are both fine-tuned on 108 datasets, indicating their role in multilingual and multi-task similarity-based retrieval applications. The ModernBERT-base-embed and Llama-3.2-3B-Instruct families show strong dataset diversity as well, with 88 and 87 datasets, respectively, across fine-tuning and quantized variants (e.g., GGUF format). Interestingly, most of these models are fine-tuned, specifically, 8 out of the top 10, highlighting a trend where base models are transformed into task-specific variants through diverse training datasets. This pattern suggests that fine-tuning remains a dominant strategy for adapting base models to downstream tasks across heterogeneous data sources.
Finding #5: Models and datasets exhibit strong bidirectional interdependence, with datasets like Mistral-7B-v0.1 spawning hundreds of models, while models such as DeBERTa-ST-AllLayers-v3.1 leverage diverse datasets to enhance adaptability, highlighting the critical role of dataset-model interactions in advancing AI.
4.6. RQ #6: Dynamic Update Evaluation
@@
Figure 5. The number of changed models and datasets on Hugging Face from June 25 to July 15, 2025.
This research question aims to understand the dynamic update of the LLM supply chain. Following the method discussed in Section 3.4, we perform a daily-based data collection by capturing how many nodes and edges are added and deleted each day.
Figure 5 illustrates the sum of daily added and deleted of six key node categories (base, fine-tuned, adapters, quantized, merge variants, and datasets) from June 25 to July 15, 2025. We make three interesting observations. (i) The daily dynamic update is significant. That is, an average of 4,622 models are changing every day, including $\sim$3,843 model additions and $\sim$779 deletions. In addition, about 1,538 datasets are changing each day, containing $\sim$1,305 dataset additions and $\sim$233 deletions. An addition occurs when a new model or dataset is uploaded to the Hugging Face platform. This includes base models, task-specific variants, and new training datasets. A deletion refers to the removal of such nodes, often due to licensing issues, privacy concerns, or contributor decisions, such as replacing outdated models or withdrawing low-quality or sensitive datasets.
(ii) Fine-tuned models dominate daily activity, averaging over 2,988 changes per day, followed by consistent contributions from adapters ($\sim$1,224/day) and datasets ($\sim$1,539/day). Noticeable spikes, such as on July 7 and July 9, align with major events like the Mistral-Fusion-v3 fine-tuning wave and dataset updates such as HFTime2025-News. (iii) Furthermore, adapter uploads peaked at 1,249 on June 28, while quantized variants reached 381 on July 9, driven by releases such as QWin-GGUF-7B. These patterns demonstrate HuggingGraph’s ability to capture evolving supply chain dynamics at a fine-grained level.
Finding #6: The LLM supply chain exhibits continuous and high-volume daily changes, driven by frequent additions and deletions of models and datasets. This reflects a rapidly evolving and highly dynamic ecosystem shaped by active contributor behavior.
4.7. RQ #7: Generability to Other AI Platforms
To validate HuggingGraph’s generalizability beyond Hugging Face, we applied our pipeline to another AI platform, Kaggle (Bojer and Meldgaard, 2021). As of July 25, 2025, Kaggle hosts 470 base and task-specific models with 3,146 variants and approximately 502K datasets. Using Kaggle’s kernel and dataset APIs, we collected 2,640 models and 105,867 datasets. This significant gap is due to lack of the models’ and datasets’ metadata. Of the datasets retrieved ($\sim$105K), only 137 datasets were included in our graph, as most lacked standardized documentation or traceable links to models. Many are standalone, poorly described, or lack contextual information, a challenge also observed on Hugging Face.
We follow the same way to construct a heterogeneous graph consisting of 2,777 nodes, which include 2,640 model nodes, comprising 59 base models, 2,410 fine-tuned, 171 quantization models, and 137 datasets. The graph contains a total of 3,990 edges, on which we observed seven types of edges, (i) base model $\rightarrow$ fine-tuned model (467 edges), (i) base model $\rightarrow$ quantization model (62 edges), (iii) fine-tuned model $\rightarrow$ fine-tuned model (1,696 edges), (iv) fine-tuned model $\rightarrow$ quantized model (107 edges), (v) quantized model $\rightarrow$ quantized model (1 edges), (vi) dataset $\rightarrow$ fine-tuned model (1,614 edges), and (vii) dataset $\rightarrow$ quantization model (43 edges).
We observed that the average degree is 1.44, indicating that the graph is sparse. We made two interesting observations. (i) The degree distribution is heavy-tailed and skewed: out of 2,777 total nodes, 2,305 nodes ($\approx$83%) have a total degree of 1. Most nodes have low degrees, while a few highly connected hubs dominate the graph. For example, in-degrees range from 0 to 4 (with NaN peaking at 4), and tensorflow/mobilenet-v1 has the highest out-degree of 64, followed by google/nnlm with 56. (ii) Furthermore, the graph contains 448 weakly connected components (WCCs), reflecting high fragmentation. However, the largest WCC includes 65 nodes, suggesting the presence of a moderately sized core subgraph.
Finding #7: The resulting graph exhibits structural properties consistent with our Hugging Face analysis, including a heavy-tailed degree distribution, sparse connectivity, and strong modular fragmentation, demonstrating the robustness and generalizability of our pipeline across platforms despite metadata limitations.
5. Use Case
Use case #1: Tracing lineage and dependencies in the LLM supply chain. In the LLM ecosystem, models are frequently built upon others through fine-tuning, adapter training, or quantization, forming complex chains of dependencies. However, when these relationships are not explicitly visible, it becomes difficult to verify where a model comes from, whether it inherits bias from upstream datasets, or if it complies with licensing constraints. HuggingGraph can be used to address this challenge by constructing the supply chain of models and datasets, uncovering both direct and derived dependencies, even when they are not formally documented. For example, it can trace how the model Meta-llama indirectly relies on a dataset like Wikimedia via Chatgpt-prompt (Figure 2). This transparency supports developers, auditors, and policymakers in validating provenance, detecting risks, and enabling trustworthy AI.
Use case #2: Identifying critical nodes and structural vulnerabilities. In the LLM ecosystem, certain models (e.g., gemma-2b) and datasets (e.g., The Pile) are reused so frequently that they become critical structural hubs, where failure or removal of them could disrupt numerous downstream dependencies. These hidden single points of failure are difficult to detect without a comprehensive view of resource interconnections. HuggingGraph can be used to address this by modeling the supply chain as a graph and analyzing node connectivity to surface highly reused models and datasets with significant inbound or outbound links. This visibility enables maintainers to safeguard vital assets and helps developers mitigate the risk of overreliance on fragile or under-maintained components.
@@ 6. Related Work
LLM supply chain perspectives in AI. LLMs mark the revolutionary era in AI, with the edge passing through an exponential growth phase. A recent paper presents LLMs key components, including model infrastructure, lifecycle, and downstream applications (Wang et al., 2024). Another work showed that reusing these models has become widespread, which encourages the sharing and adaptation of base models on a larger scale (Jiang et al., 2023). An open-source AI ecosystem, such as Hugging Face, hosts a broad range of LLMs and datasets and, therefore, plays an important role in democratizing AI technologies (Riva et al., 2024). The base models are key ingredients in such an ecosystem (Xu et al., 2024b). They capture vast amounts of information from various datasets, allowing for effective task-specific variants like fine-tuning (Han et al., 2021). This trend underscores the significant impact of technological advancements on AI democratization and innovation (Cupać et al., 2024).
@@
Relationship analysis between LLM models and datasets. A recent study investigates the practical adaptation of base models to specific areas and tasks. Multitask fine-tuning has demonstrated the potential to enhance performance on target tasks with scarce labels (Xu et al., 2024a). In plant phenotyping, adapting vision-based models by techniques like adapter tuning and decoder tuning has shown results comparable to those of leading task-specific models (Chen et al., 2023). The Quadapter technique for language models tackles quantization difficulties by incorporating learnable parameters that scale activations channel-wise, mitigating overfitting during quantization-aware training (Park et al., 2022). These findings underscore the efficacy of various adaptation approaches in improving base model performance across areas.
@@
From the above, we can see that models heavily depend on each other for fine-tuning, adaptation, quantization or merger. Building on these insights, two other works have indicated that serious security vulnerabilities may encumber the LLM supply chain to a few developers (Kathikar et al., 2023; Osborne et al., 2024). These factors may affect the diversity of innovation. The authors relate several challenges arising during the software engineering, security, and privacy of different relations and components involved in model creation and deployment (Huang et al., 2023; Wang et al., 2024). Moreover, the vulnerabilities can be transferable from one model to another model during the fine-tuning process (Hajipour et al., 2024).
@@ 7. Conclusion
This project studies the relationships between models and datasets in the LLM ecosystem, which are the central parts of the LLM supply chain. First, we design a methodology to systematically collect the supply chain information of LLMs. With that, we construct a large directed heterogeneous graph, having 402,654 nodes and 453,469 edges. Then, we perform different types of analysis and make multiple interesting findings.